OpenStudy (anonymous):

4^x=12 ....solving exponential equations with equal bases?

sam (.sam.):

4^(x)=12 ln(4^(x))=ln(12) xln(4)=ln(12) x=(ln(12))/(ln(4)) x=1.7925

OpenStudy (anonymous):

options are: a)1.79 b)19.93 c)0.27 d)0.56

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank u guys so much!!!